OpenStudy (anonymous):

hi,i want to ask 1 question the question is like this if a and b are positive numbers,find the maximum value of f(x)=x^a(1-x)^b where x is more and equal to 0 and x is less and equal to 1.

zepdrix (zepdrix):

Hi, \(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

Mmm so what are we doing? Using basic calc methods to figure this out?

zepdrix (zepdrix):

\[\Large\rm f(x)=x^a(1-x)^b,\qquad 0\le x\le 1\]

zepdrix (zepdrix):

\[\Large\rm a>0, \qquad b>0\]

OpenStudy (anonymous):

it is ok if i just assume any positive number for a and b?

zepdrix (zepdrix):

Well your final answer will probably involve a and b, so you don't want to plug in values to try and simplify the problem, I think that might end up causing some trouble. If that's what you were asking..

zepdrix (zepdrix):

Applying product rule to take our derivative: \[\Large\rm f'(x)=a x^{a-1}(1-x)^b-b x^a(1-x)^{b-1}\]That step make sense?

OpenStudy (anonymous):

i have try product rule and i've got like this ax^(a−1)(1−x)^b=x^ab(1−x)^(b−1) ax^(a−1)x^(−a)(1−x)^b=b(1−x)^(b−1) ax^−1(1−x)^b=b(1−x)^(b−1) a=xb(1−x)^(b−1)(1−x)^−b a=xb(1−x)^−1 a/b=x/1−x a−ax=bx a=bx+ax a=x(b+a) x=a/(b+a)